The rolling resistance of steel is typically mu_r
Answer #1

The frictional force will be:

Ff = 160,000 x 9.8 x 0.002 = 3136 N

acceleration would be:

a = Ff/m = 3136/160000 = 0.0196 m/s^2

we know that, a = v/t => t = v/a

t = 12/0.0196 = 612.2 sec

The distance travelled will be:

D = Vavg x t = 12/2 x 612.2 = 3673.5 m

Hence, t = 612.2 s and D = 3673.5 m.